OpenStudy (anonymous):

find the general solution.. 1. y''-2y'-2y=0 2. y''''-8y''-9y=0 3. y''''+13y''+36y=0 4. y'''''-15y''''+84y'''-220y''-275y'=0 5. y'''''+12y''''+104y'''+408y''-1156y'=0

OpenStudy (anonymous):

1- D2-2D-2=0 (D-1)^2=3 D=1+-sqrt(3) y=c1exp(1+sq3)x+c2exp(1-sq3)x